the perimemter of a rectangle = 150 yds, and the length is 2 times the width. What is the length and width?
Answer by checkley77(12844)

L=2W
2L+2W=150
2*2W+2W=150
4W+2W=150
6W=150
W=150/6
W=25 ANS. FOR THE WIDTH.
L=2*25=50 ANS. FOR THE LENGTH.
PROOF:
2*50+2*25=150
100+50=150
150=150